the product of two consecutive positive integers is 11 more than their sum. find the intergers
Answer by mananth(16949)

let the integers be x & x+1
Product = x(x+1)
...
x+x+1+11
...
x(x+1)=x+x+1+11
x^2+x=2x+12
-2x -2x
x^2-2x+x=12
x^2-x=12
-12 -12
x^2-x-12=0
(x-4)(x+3)=12
x= 4 OR -3
...
4,5 OR -3,-2
...
m.ananth@hotmail.ca